OpenStudy (anonymous):

Which statement is true about the product square root of 2(3square root of 2 + square root of 14)? It is rational and equal to 12. It is rational and equal to 18. It is irrational and equal to 6 + 2square root of 18. It is irrational and equal to 3 +square root of 18.

OpenStudy (anonymous):

Which statement is true about the product square root of \[\sqrt{2}(3\sqrt{2}+\sqrt{18})\] It is rational and equal to 12. It is rational and equal to 18. It is irrational and equal to 6 + 2square root of 18. It is irrational and equal to 3 +square root of 18.

OpenStudy (anonymous):